OpenStudy (aroub):

At 9:00 A.M. a car leaves a certain town at a speed of 90 km/hr. After 60 minutes another car leaves the same town to follow the first car. The two cars meet at 12:00 P.M. a) Find the distance travelled by each of the two cars. b) Find the speed of the second car. ( & it's a motion problem btw. )

OpenStudy (anonymous):

"motion problem" Aha!, En passant we here call it "Time and distance" problem :)

OpenStudy (anonymous):

Okay, after one hour how much distance did the first car had traveled ?

OpenStudy (anonymous):

Okay if car travel 90km/hr how much does it travel in 60 mints ?

OpenStudy (aroub):

d=r*t d=90*60 d=5400 right?

OpenStudy (anonymous):

What is d=r*t ?

OpenStudy (ksaimouli):

change in to min when u get the distance

OpenStudy (ksaimouli):

speed=distance over time

OpenStudy (ksaimouli):

u would get it

OpenStudy (anonymous):

Don't think of formulas... just think simply .. what do we mean when we say " My speed is 60km/hr?"

OpenStudy (aroub):

distance = rate ( or speed) times time ---> formula you use it in 'motion problems' or 'Time and distance problems'

OpenStudy (hoblos):

a) the first car traveled 3 hours at a speed of 90 km/hr => it traveled 90*3= 270Km and since they met by this time and started from the same place then the second car also traveled the same distance 270Km b)the second car left after 60min => at 10:00AM => it traveled for 2 hrs its speed = distance/time = 270km/2hrs = 135km/hr

OpenStudy (anonymous):

three hours times 90 kilometers per hours is 270 kilometers right?

OpenStudy (anonymous):

@hoblos: nice answer, but she don't understand the concept of speed.

OpenStudy (anonymous):

i guess if they meet at noon the distance traveled is the same for both cars

OpenStudy (anonymous):

of course maybe it is noon on the next day.

OpenStudy (aroub):

Yes the distance is the same

OpenStudy (anonymous):

saifoo! happy new year!

OpenStudy (saifoo.khan):

Happy new year Sat Sir. :D

OpenStudy (aroub):

Okay I get it now.. Thanks Hoblos :D

OpenStudy (anonymous):

sir? lol

OpenStudy (saifoo.khan):

Yeah! =D

OpenStudy (anonymous):

back atcha

OpenStudy (saifoo.khan):

Lol, c'mon!